In February 2015 Treasury 4 5/8s of 2042 offered a semiannually compounded yield to maturity of 2.84%. Recognizing that coupons are paid semiannually, calculate the bond's price. Assume face value is $1,000. (Do not round intermediate calculations. Round your answer to 2 decimal places.)

Bond price            $
Answer #1

Since Treasury 4 5/8s is a Zero Coupon Bond, it implies coupons are not provided

Bond's Price = INT(1-(1+r)-n)/r+A(1+r)-n

Coupon not provided therefore, INT(1-(1+r)-n)/r = 0

n = (2042-2015)= 27 years

=0+A(1+r)-n

=0+1000(1+.0284)-27

=$469.49